# SG-GS: Photo-realistic Animatable Human Avatars with Semantically-Guided Gaussian Splatting

### v1

@@ Abstract
Reconstructing photo-realistic animatable human avatars from monocular videos remains challenging in computer vision and graphics. Recently, methods using 3D Gaussians to represent the human body have emerged, offering faster optimization and real-time rendering. However, due to ignoring the crucial role of human body semantic information which represents the intrinsic structure and connections within the human body, they fail to achieve fine-detail reconstruction of dynamic human avatars. To address this issue, we propose SG-GS, which uses semantics-embedded 3D Gaussians, skeleton-driven rigid deformation, and non-rigid cloth dynamics deformation to create photo-realistic animatable human avatars from monocular videos. We then design a Semantic Human-Body Annotator (SHA) which utilizes SMPL’s semantic prior for efficient body part semantic labeling. The generated labels are used to guide the optimization of Gaussian semantic attributes. To address the limited receptive field of point-level MLPs for local features, we also propose a 3D network that integrates geometric and semantic associations for human avatar deformation. We further implement three key strategies to enhance the semantic accuracy of 3D Gaussians and rendering quality: semantic projection with 2D regularization, semantic-guided density regularization and semantic-aware regularization with neighborhood consistency. Extensive experiments demonstrate that SG-GS achieves state-of-the-art geometry and appearance reconstruction performance. The code will be made available.
@@
footnotetext: Haoyu Zhao completed this work during an internship at Shanghai Jiao Tong University.
Introduction
@@
Creating photo-realistic, animatable human avatars from monocular videos has substantial implications for gaming (Zackariasson and Wilson 2012), extended reality storytelling (Healey et al. 2021), and tele-presentation (Ho et al. 2023). In this work, we are dedicated to create high-quality photo-realistic, animatable human avatars from monocular videos with semantics embedded 3D Gaussians.
@@
Recent advances in implicit neural fields (Mildenhall et al. 2021; Oechsle, Peng, and Geiger 2021; Niemeyer et al. 2020; Wang et al. 2022a; Sitzmann et al. 2021) enable high-quality reconstruction of geometry (Xu, Alldieck, and Sminchisescu 2021; Wang et al. 2022b; Guo et al. 2023) and appearance (Jiang et al. 2022; Li et al. 2022; Raj et al. 2021; Yu et al. 2023; Weng et al. 2022) of clothed human bodies from sparse multi-view or monocular videos. However, they often employ large MLPs, which makes training and rendering computationally demanding and inefficient.
@@
Point-based rendering (Rückert, Franke, and Stamminger 2022; Su, Bagautdinov, and Rhodin 2023; Zheng et al. 2023) has emerged as an efficient alternative to NeRFs, offering significantly faster rendering. The recently proposed 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) achieves state-of-the-art novel view synthesis performance on static scenes with significantly reduced inference time and faster training. 3DGS has inspired several recent works in animatable human avatar creation (Lei et al. 2024; Moreau et al. 2024; Shao et al. 2024; Hu et al. 2024b; Kocabas et al. 2024; Wang et al. 2024; Qian et al. 2024; Hu et al. 2024a). However, these methods often overlook crucial semantic information that represents the connections within the human body, leading to issues in maintaining anatomical coherence during motion and preserving fine details such as muscle definition and skin folds in various poses.
@@
To this end, we propose SG-GS, a Semantically-Guided 3D human model using Gaussian Splatting representation. Our SG-GS first integrates a skeleton-driven rigid deformation, and a non-rigid cloth dynamics deformation to coordinate the movements of individual Gaussians during animation. We then introduce a Semantic Human-Body Annotator (SHA), which leverages SMPL’s (Loper et al. 2015) human semantic prior for efficient body part semantic labeling. These part labels are used to guide the optimization of 3D Gaussian’s semantic attribute. Point-level MLPs have a limited receptive field for capturing local geometric and semantic features. To address this, we propose a 3D geometry- and semantic-aware network to learn body geometric and semantic associations and integrate them into the avatar deformation. We further implement three key strategies to enhance semantic accuracy of 3D Gaussians and rendering quality: semantic projection with 2D regularization, semantic-guided density regularization and semantic-aware regularization with neighborhood consistency. Our experimental results demonstrate that SG-GS achieves superior performance compared to current state-of-the-art approaches (Qian et al. 2024; Hu et al. 2024b, a) in animatable avatar creation from monocular inputs.
In summary, our work makes the following contributions:
@@
  • We propose a 3D geometry- and semantic-aware network to capture detailed local geometric and semantic information about the human body.
@@
  • We introduce semantic projection with 2D regularization, semantic-aware regularization with neighborhood consistency, and semantic-guided density regularization to improve semantic accuracy and rendering quality.
Related Work
Neural Rendering for Human Avatars
@@
Since the introduction of Neural Radiance Fields (NeRF) (Mildenhall et al. 2021), there has been a surge of research on neural rendering for human avatars (Li et al. 2022; Liu et al. 2021; Li et al. 2023; Peng et al. 2021b; Wang et al. 2022b). Though, NeRF is designed for static objects, HumanNeRF (Weng et al. 2022) extend the NeRF to enable capturing a dynamic moving human using just a single monocular video. Neural Body (Peng et al. 2021b) associates a latent code to each SMPL (Loper et al. 2015) vertex to encode the appearance, which is transformed into observation space based on the human pose. Furthermore, Neural Actor (Liu et al. 2021) learns a deformable radiance field with SMPL (Loper et al. 2015) as guidance and utilizes a texture map to improve its final rendering quality. Posevocab (Li et al. 2023) designs joint-structured pose embeddings to encode dynamic appearances under different key poses, enabling more effective learning of joint-related appearances. However, a major limitation of NeRF-based methods is that NeRFs are slow to train and render.
@@
Some works focus on achieving fast inference and training times for NeRF models of human avatars, including approaches that use explicit representations such as learning a function at grid points (Chen et al. 2022; Reiser et al. 2021), using hash encoding (Müller et al. 2022), or altogether discarding the learnable component (Fridovich-Keil et al. 2022). iNGP (Müller et al. 2022) uses the underlying representation for articulated NeRFs, and enable interactive rendering speeds (15 FPS). (Chen et al. 2023) generates a pose-dependent UV volume, but its UV volume generation is not fast (20 FPS). In contrast to all these works, SG-GS achieves state-of-the-art rendering quality and speed (25 FPS) with less training time from monocular video inputs.
Dynamic 3D Gaussians for Human Avatars
@@
Point-based rendering (Rückert, Franke, and Stamminger 2022; Su, Bagautdinov, and Rhodin 2023; Zheng et al. 2023) has proven to be an efficient alternative to NeRFs for fast inference and training. Extending point clouds to 3D Gaussians, 3D Gaussian Splatting (3DGS) (Kerbl et al. 2023) models the rendering process by splatting a set of 3D Gaussians onto the image plane via alpha blending. This approach achieves state-of-the-art rendering quality with real-time inference speed for novel views and fast training.
@@
Given the impressive performance of 3DGS in both quality and speed, numerous works have further explored the 3D Gaussian representation for dynamic scene reconstruction. D-3DGS (Luiten et al. 2024) is proposed as the first attempt to adapt 3DGS into a dynamic setup. Other works (Wu et al. 2024; Yang et al. 2024b; Zhao et al. 2024) model 3D Gaussian motions with a compact network or 4D primitives, resulting in highly efficient training and real-time rendering.
@@
The application of 3DGS in dynamic 3D human avatar reconstruction is just beginning to unfold (Jiang et al. 2024; Lei et al. 2024; Qian et al. 2024; Hu et al. 2024b; Kocabas et al. 2024). Human Gaussian Splatting (Moreau et al. 2024) showcase 3DGS as an efficient alternative to NeRF. Splattingavatar (Shao et al. 2024) and Gomavatar (Wen et al. 2024) extends lifted optimization to simultaneously optimize the parameters of the Gaussians while walking on the triangle mesh. While these methods have made significant progress, they often overlook the crucial role of semantic information. It is a key focus of our SG-GS.
Preliminaries
@@
SMPL (Loper et al. 2015). The SMPL model is a widely-used parametric 3D human body model that efficiently represents body shape and pose variations. In our work, We utilize SMPL’s Linear Blend Skinning (LBS) algorithm to transform points from canonical space to observation space, enabling accurate body deformation across different poses. We also leverage SMPL’s body priors for semantic information, enhancing our model’s understanding of body structure and part relationships. This semantic information is crucial for improving the quality and consistency of our human avatar reconstruction.
@@
3D Gaussian Splatting (3DGS) (Kerbl et al. 2023). 3DGS explicitly represents scenes using point clouds, where each point is modeled as a 3D Gaussian defined by a covariance matrix $\Sigma$ and a center point $\mathcal{X}$, the latter referred to as the mean. The value at point $\mathcal{X}$ is:
$G(\mathcal{X})=e^{-\frac{1}{2}\mathcal{X}^{T}\Sigma^{-1}\mathcal{X}}.$ (1)
@@
For differentiable optimization, the covariance matrix $\Sigma$ is decomposed into a scaling matrix $\mathcal{S}$ and a rotation matrix $\mathcal{R}$, such that $\Sigma=\mathcal{R}\mathcal{S}\mathcal{S}^{T}\mathcal{R}^{T}$. In practice, $\mathcal{S}$ and $\mathcal{R}$ are also represented by the diagonal vector $s\in\mathbb{R}^{N\times 3}$ and a quaternion vector $r\in\mathbb{R}^{N\times 4}$, respectively. In rendering novel views, differential splatting, as introduced by (Yifan et al. 2019) and (Zwicker et al. 2001), involves applying a viewing transformation $W$ along with the Jacobian matrix $J$ of the affine approximation of the projective transformation. This process computes the transformed covariance matrix as: $\Sigma^{\prime}=JW\Sigma W^{T}J^{T}$. The color and opacity at each pixel are computed from the Gaussian’s representation $G(\mathcal{X})=e^{-\frac{1}{2}\mathcal{X}^{T}\Sigma^{-1}\mathcal{X}}$. The blending of $N$ ordered points overlapping a pixel is given by the formula:
@@
Figure 1: Our framework for creating photo-realistic animatable avatars from monocular videos. We initialize a set of 3D Gaussians in the canonical space by sampling 6,890 points from the SMPL model and assign the semantic attributes of Gaussians to each point. We first integrate a skeleton-driven rigid deformation and a non-rigid cloth dynamics deformation to deform human avatars from canonical space $\mathcal{G}_{c}$ to observation space $\mathcal{G}_{o}$. Then, we introduce a Semantic Human-Body Annotator (SHA), which leverages SMPL’s human body semantic prior for efficient semantic labeling. These labels are used to guide the optimization of 3D Gaussian’s semantic attribute $\mathcal{O}$. To utilize 3D geometric and semantic information effectively, we also propose a 3D geometry- and semantic-aware network to learn body geometric and semantic associations and integrate them into learning the 3D deformation. To enhance semantic accuracy and render quality, we implement semantic projection with 2D regularization, semantic-guided density regularization and semantic-aware regularization with neighborhood consistency.
Method
In this section, we illustrate the pipeline of our SG-GS in Fig. 1. The inputs to our method include images $X=\{x_{i}\}^{N}_{i=1}$ obtained from monocular videos, fitted SMPL parameters $P=\{p_{i}\}^{N}_{i=1}$, and paired foreground masks $M=\{m_{i}\}^{N}_{i=1}$ of images. SG-GS optimizes 3D Gaussians in canonical space, which are then deformed to match the observation space and rendered from the provided camera view. To integrate semantic information about body parts into the 3D Gaussian optimization process, we divide the human body into 5 distinct parts and represent the labels using one-hot encoding, stored as semantic attribute $\mathcal{O}\in\mathbb{R}^{10}$.
Non-rigid and Rigid Deformation
Inspired by (Weng et al. 2022; Qian et al. 2024), We decompose human deformation into two key components: 1) a non-rigid element capturing pose-dependent cloth dynamics, and 2) a rigid transformation governed by the human skeletal structure.
We employ a non-rigid deformation network, that takes the canonical position $\mathcal{X}_{c}$ of the 3D Gaussians ${\mathcal{G}_{c}}$ in canonical space and a pose latent code as input. This pose latent code encodes SMPL parameters $p_{i}$ using a lightweight hierarchical pose encoder (Mihajlovic et al. 2021). The network then outputs offsets for various parameters of ${\mathcal{G}_{c}}$:
@@
Semantic Human-Body Annotator
Most current animatable human avatar creation methods just use SMPL (Loper et al. 2015) model for its pose-aware shape priors, neglecting its inherent semantic information. We argue that accurate semantic information can improve rendering quality, and we will further demonstrate this in the experimental section.
@@
where $\mathcal{SHA}$ denotes our Semantic Human-Body Annotator. We use the generated human body semantic labels $m_{s}$ to ensure the accuracy of Gaussian’s semantic attribute. (described in semantic projection with 2D regularization).
Geometric and Semantic Feature Learning
Maintaining motion consistency across local neighborhoods is essential in modeling non-rigid deformation. However, point-level MLPs are limited by a small receptive field, which restricts their capability to capture the local geometric and semantic features of point clouds. To address this, we propose a 3D geometry- and semantic-aware network that effectively captures the human body’s local geometric and semantic structure in canonical space.
We employ sparse convolution (Graham and Van der Maaten 2017) on sparse voxels to extract local geometric and semantic features across varying receptive fields, following the method outlined in (Lu et al. 2024). Given the position $\mathcal{X}_{c}$ of the Gaussians $\mathcal{G}_{c}$ as a point cloud, we initially convert it into voxels by partitioning the space using a fixed grid size $v$.

### v2

# SG-GS: Topology-aware Human Avatars with Semantically-guided Gaussian Splatting

@@ Abstract
Reconstructing photo-realistic and topology-aware animatable human avatars from monocular videos remains challenging in computer vision and graphics. Recently, methods using 3D Gaussians to represent the human body have emerged, offering faster optimization and real-time rendering. However, due to ignoring the crucial role of human body semantic information which represents the explicit topological and intrinsic structure within human body, they fail to achieve fine-detail reconstruction of human avatars. To address this issue, we propose SG-GS, which uses semantics-embedded 3D Gaussians, skeleton-driven rigid deformation, and non-rigid cloth dynamics deformation to create photo-realistic human avatars. We then design a Semantic Human-Body Annotator (SHA) which utilizes SMPL’s semantic prior for efficient body part semantic labeling. The generated labels are used to guide the optimization of semantic attributes of Gaussian. To capture the explicit topological structure of the human body, we employ a 3D network that integrates both topological and geometric associations for human avatar deformation. We further implement three key strategies to enhance the semantic accuracy of 3D Gaussians and rendering quality: semantic projection with 2D regularization, semantic-guided density regularization and semantic-aware regularization with neighborhood consistency. Extensive experiments demonstrate that SG-GS achieves state-of-the-art geometry and appearance reconstruction performance. Our project is at https://sggs-projectpage.github.io/.
Figure 1: We propose an efficient method for creating topology-aware human avatars from just videos, ensuring both photo-realistic human appearance and accurate anatomical structure. Our method achieve better quality to the most recent state-of-the-art methods [39, 10, 32].
@@
footnotetext: * Equal contributions.footnotetext: †Corresponding Author.footnotetext: Haoyu Zhao completed this work during an internship at Shanghai Jiao Tong University.
1 Introduction
@@
Creating photo-realistic human avatars from monocular videos has immense potential value in industries such as gaming [47], extended reality storytelling [7], and tele-presentation [8]. In this work, we are dedicated to create high-quality photo-realistic human avatars from monocular videos with semantics embedded 3D Gaussians.
@@
Recent advances in implicit neural fields [26, 36] enable high-quality reconstruction of geometry [42, wang2022arah, 6] and appearance [13, 19, 46, 40] of clothed human bodies from sparse multi-view or monocular videos. However, they often employ large MLPs, which makes training and rendering computationally demanding and inefficient.
@@
Point-based rendering [49] has emerged as an efficient alternative to NeRFs, offering significantly faster rendering speed. The recently proposed 3D Gaussian Splatting (3DGS) [15] achieves state-of-the-art novel view synthesis performance with significantly reduced inference time and faster training. 3DGS has inspired several recent works in human avatar creation [17, 27, 35, 10, 16, 37, 32, 9]. However, these methods often overlook crucial semantic information that represents the explicit topological structure within the human body, leading to issues in maintaining anatomical coherence during motion and preserving fine details such as muscle definition and skin folds in various poses.
@@
To this end, we propose SG-GS, a Semantically-Guided 3D human model using Gaussian Splatting representation, as shown in Fig. 1. SG-GS first integrates a skeleton-driven rigid deformation, and a non-rigid cloth dynamics deformation to coordinate the movements of individual Gaussians during animation. We then introduce a Semantic Human-Body Annotator (SHA), which leverages SMPL’s [22] human semantic prior for efficient body part semantic labeling. These part labels are used to guide the optimization of 3D Gaussian’s semantic attribute. To learn topological relationships between human body parts, we propose a 3D topology- and geometry-aware network to learn body geometric and topological associations and integrate them into the avatar deformation. We further implement three key strategies to enhance semantic accuracy of 3D Gaussians and rendering quality: semantic projection with 2D regularization, semantic-guided density regularization and semantic-aware regularization with neighborhood consistency. Our experimental results demonstrate that SG-GS achieves superior performance compared to current SOTA approaches in avatar creation from monocular inputs. In summary, our work makes the following contributions:
@@
  • We propose a 3D topology and geometry-aware network to capture topology and geometry information within the human body.
@@
  • We introduce semantic projection with 2D regularization, semantic neighborhood-consistent regularization, and semantic-guided density regularization to enhance semantic accuracy and rendering quality.
2 Related Work
2.1 Neural Rendering for Human Avatars
@@
Since the introduction of Neural Radiance Fields (NeRF) [26], there has been a surge of research on neural rendering for human avatars [19, 21, 20, 31, wang2022arah]. Though, NeRF is designed for static objects, HumanNeRF [40] extend the NeRF to enable capturing a dynamic moving human using just a single monocular video. Neural Body [31] associates a latent code to each SMPL [22] vertex to encode the appearance, which is transformed into observation space based on the human pose. Furthermore, Neural Actor [21] learns a deformable radiance field with SMPL [22] as guidance and utilizes a texture map to improve its final rendering quality. Posevocab [20] designs joint-structured pose embeddings to encode dynamic appearances under different key poses, enabling more effective learning of joint-related appearances. However, a major limitation of NeRF-based methods is that NeRFs are slow to train and render.
@@
Some works focus on achieving fast inference and training times for NeRF models of human avatars, including approaches that use explicit representations such as learning a function at grid points [1], using hash encoding [28], or altogether discarding the learnable component [3]. iNGP [28] uses the underlying representation for articulated NeRFs, and enable interactive rendering speeds (15 FPS). [2] generates a pose-dependent UV volume, but its UV volume generation is not fast (20 FPS). In contrast to all these works, SG-GS achieves state-of-the-art rendering quality and speed (25 FPS) with less training time.
2.2 Dynamic 3D Gaussians for Human Avatars
@@
Point-based rendering [34, 49] has proven to be an efficient alternative to NeRFs for fast inference and training. Extending point clouds to 3D Gaussians, 3D Gaussian Splatting (3DGS) [15] models the rendering process by splatting a set of 3D Gaussians onto the image plane via alpha blending. This approach achieves SOTA rendering quality with fast inference speed for novel views.
@@
Given the impressive performance of 3DGS in both quality and speed, numerous works have further explored the 3D Gaussian representation for dynamic scene reconstruction. D-3DGS [24] is proposed as the first attempt to adapt 3DGS into a dynamic setup. Other works [41, 44, 48] model 3D Gaussian motions with a compact network or 4D primitives, resulting in highly efficient training and real-time rendering.
@@
The application of 3DGS in dynamic 3D human avatar reconstruction is just beginning to unfold [14, 17, 32, 10, 16]. Human Gaussian Splatting [27] showcase 3DGS as an efficient alternative to NeRF. Splattingavatar [35] and Gomavatar [39] extends lifted optimization to simultaneously optimize the parameters of the Gaussians while walking on the triangle mesh. While these methods have made significant progress, they often overlook the crucial role of semantic information which is related to topological relationships between human body parts. It is a key focus of our SG-GS.
3 Preliminaries
@@
SMPL [22]. The SMPL model is a widely-used parametric 3D human body model that efficiently represents body shape and pose variations. In our work, We utilize SMPL’s Linear Blend Skinning (LBS) algorithm to transform points from canonical space to observation space, enabling accurate body deformation across different poses. We also leverage SMPL’s body priors to enhance the model’s understanding of body structure, improving the quality and consistency of human avatar reconstruction.
@@
3D Gaussian Splatting (3DGS) [15]. 3DGS explicitly represents scenes using point clouds, where each point is modeled as a 3D Gaussian defined by a covariance matrix $\Sigma$ and a center point $\mathcal{X}$, the latter referred to as the mean. The value at point $\mathcal{X}$ is: $G(\mathcal{X})=e^{-\frac{1}{2}\mathcal{X}^{T}\Sigma^{-1}\mathcal{X}}$.
@@
For differentiable optimization, the covariance matrix $\Sigma$ is decomposed into a scaling matrix $\mathcal{S}$ and a rotation matrix $\mathcal{R}$, such that $\Sigma=\mathcal{R}\mathcal{S}\mathcal{S}^{T}\mathcal{R}^{T}$. In practice, $\mathcal{S}$ and $\mathcal{R}$ are also represented by the diagonal vector $s\in\mathbb{R}^{N\times 3}$ and a quaternion vector $r\in\mathbb{R}^{N\times 4}$, respectively. In rendering novel views, differential splatting, as introduced by [45] and [50], involves applying a viewing transformation $W$ along with the Jacobian matrix $J$ of the affine approximation of the projective transformation. This process computes the transformed covariance matrix as: $\Sigma^{\prime}=JW\Sigma W^{T}J^{T}$. The color and opacity at each pixel are computed from the Gaussian’s representation $G(\mathcal{X})=e^{-\frac{1}{2}\mathcal{X}^{T}\Sigma^{-1}\mathcal{X}}$. The pixel color $\mathcal{C}$ is computed by blending $N$ ordered 3D Gaussian splats that overlap at the given pixel, using the formula:
@@
Figure 2: Our framework for creating photo-realistic animatable avatars from monocular videos. We initialize a set of 3D Gaussians in the canonical space by sampling 6,890 points from the SMPL model and assign the semantic attributes of Gaussians to each point. We first integrate a skeleton-driven rigid deformation and a non-rigid cloth dynamics deformation to deform human avatars from canonical space $\mathcal{G}_{c}$ to observation space $\mathcal{G}_{o}$. Then, we introduce a Semantic Human-Body Annotator (SHA), which leverages SMPL’s human body semantic prior for efficient semantic labeling. These labels are used to guide the optimization of 3D Gaussian’s semantic attribute $\mathcal{O}$. We also propose a 3D topology and geometry-aware network to learn body topological and geometric associations and integrate them into learning the 3D deformation. To enhance semantic accuracy and render quality, we implement semantic projection with 2D regularization, semantic-guided density regularization and semantic-aware regularization with neighborhood consistency.
4 Method
In this section, we illustrate the pipeline of our SG-GS in Fig. 2. The inputs to our method include images $X=\{x_{i}\}^{N}_{i=1}$ obtained from monocular videos, fitted SMPL parameters $P=\{p_{i}\}^{N}_{i=1}$, and paired foreground masks $M=\{m_{i}\}^{N}_{i=1}$ of images. SG-GS optimizes 3D Gaussians in canonical space, which are then deformed to match the observation space and rendered from the provided camera view. For a set of 3D Gaussians, we store the following properties at each point: position $\mathcal{X}\in\mathbb{R}^{3}$, color defined by spherical harmonic (SH) coefficients $\mathcal{C}\in\mathbb{R}^{k}$ (where $k$ is the number of SH functions), opacity $\alpha\in\mathbb{R}$, rotation factor $r\in\mathbb{R}^{4}$, and scaling factor $s\in\mathbb{R}^{3}$. To integrate semantic information about body parts into the 3D Gaussian optimization process and learn the topological structure of the human body, we divide the human body into 5 distinct parts, as shown in Fig. 1. We represent the labels using one-hot encoding, stored as semantic attribute $\mathcal{O}\in\mathbb{R}^{10}$.
4.1 Non-rigid and Rigid Deformation
Inspired by [40, 32], We decompose human deformation into two key components: 1) a non-rigid element capturing pose-dependent cloth dynamics, and 2) a rigid transformation governed by the human skeletal structure.
We employ a non-rigid deformation network, that takes the canonical position $\mathcal{X}_{c}$ of the 3D Gaussians ${\mathcal{G}_{c}}$ in canonical space and a pose latent code as input. This pose latent code encodes SMPL parameters $p_{i}$ using a lightweight hierarchical pose encoder [25] into $\mathcal{Z}_{p}$. The network then outputs offsets for various parameters of ${\mathcal{G}_{c}}$:
@@ 4.1 Non-rigid and Rigid Deformation
4.2 Semantic Human-Body Annotator
Most current animatable human avatar creation methods just use SMPL [22] model for its pose-aware shape priors, neglecting its inherent semantic information. We argue that semantic information contains topological relationships within human body which can improve rendering quality during complex motion deformations. We will further demonstrate this in the experimental Section. 5.3.
@@ 4.2 Semantic Human-Body Annotator
where $\mathcal{SHA}$ denotes our Semantic Human-Body Annotator. We use the generated human body semantic labels $m_{s}$ to supervise the Gaussian’s semantic attribute. (described in semantic projection with 2D regularization in Section. 4.4).
While there are pre-trained networks for human parsing, such as SCHP [18] and Graphonomy [4], they are designed to segment both clothing and human body parts jointly. In contrast, our work focuses on leveraging semantic information to learn the topological relationships between different body parts. The objectives and tasks of these networks do not fully align with our needs, limiting their ability to model the geometric structure and topological connections of the human body. Their clothing segmentation can also introduce noise, hindering accurate body topology learning.
4.3 Topological and Geometric Feature Learning
To jointly learn and embed topology and geometry information to human avatar deformation, we propose a 3D topology- and geometry-aware network that effectively captures the human body’s local topological and geometric structure in canonical space.
We treat 3D Gaussians as a point cloud. Point-level MLPs are limited by a small receptive field, which restricts their capability to capture the local geometric and topological features. Therefore, we employ sparse convolution [5] on sparse voxels to extract local topological and geometric features across varying receptive fields, following the method outlined in [23]. Given the position $\mathcal{X}_{c}$ of the Gaussians $\mathcal{G}_{c}$ as a point cloud, we initially convert it into voxels by partitioning the space using a fixed grid size $v$.